# Mapping bipartite networks into multidimensional hyperbolic spaces

### v1

@@ II.4 Embedding of real bipartite networks
Using B-Mercator, we embedded the Unicodelang dataset in various dimensions. The inferred embeddings are able to reproduce the topological properties of the network (see Supplementary Figure 15). In Figure 4, we show a dual embedding representation of this dataset in dimension $D=1$. First, we focus on the countries in which a given language is used. For instance, in Figure 4a, we plot all countries in which English is spoken, i.e., the neighbors of the English language in the bipartite network. One can notice that English is located closer to India or the Philippines than to the United States of America (USA). This can be explained by the fact that many different languages are spoken in the USA, which influence its position in the bipartite map. Indeed, in Figure 4b, where we plot all neighbors of the Spanish language, the USA is located close to Spanish language. Lastly, in Figure 4c one can observe that countries from the French colonialism are concentrated in the similarity space and lie close to the French language, in contrast to France and other European countries. We can also shift the perspective: instead of examining the neighbors of each language, we can analyze the neighbors of a given country. In the bottom row of Figure 4, we plot the language neighbors of India, China, and Brazil. Countries are often located in the embedding space close to the most widely spoken language. Additional examples are provided in Supplementary Figure 13, where we explore the neighbors of languages such as Hindi, Swahili, Catalan, Persian, Korean, Dutch, Russian, and Arabic. Similarly, in Supplementary Figure 14, we depict the neighbors of countries including Canada, Turkey, Indonesia, Cameroon, Tanzania, Greece, the Philippines, and Bolivia.
@@
Figure 4: Visualization of the bipartite-$\mathbb{S}^{1}$ embedding of the Unicodelang dataset per country or language. Panels (a, b, c) show countries where a given language is spoken, i.e., the neighbours of the language node. The size of the nodes is proportional to the number of language speakers in that country. The color corresponds to the geographical region in which the country is located. A star marker indicates the position of a given language. In panels (d, e, f), we depict all languages spoken in a given country, i.e., the neighbours of the country node. The size of the nodes is proportional to the fraction of speakers of a given language. The color represents that language’s script. A cross marker indicates the position of a given country.
The hyperbolic bipartite embedding enables us to examine the concentration of countries sharing the same language. We select the top 15 languages with the highest degree and compute the angular distance from each language to its neighboring countries. A small average angular distance may indicate that these countries are highly similar, whereas a broad distribution suggests a more international language. Figure 5a displays the diversity for these high-degree languages. As expected, the angular distance distribution for English is broad, indicating connections with countries distributed throughout the similarity space. Interestingly, the angular distances for Fula are relatively small. Fula, a Senegambian language spoken primarily in West and Central Africa, is concentrated in a specific region of our embedding space. Similarly, we can examine the linguistic diversity within each country. Following the previous approach, we selected the top 15 countries with the highest degree and computed the angular distances to each neighboring language. In a given country, if the languages are more similar, their angular positions in the similarity space should be more concentrated. Figure 5b shows that India and Canada exhibit a broader distribution of angular distances, reflecting the presence of a diverse array of languages. In contrast, Russia and Cameroon display a narrower distribution, indicating a more homogeneous set of languages. Notably, all these countries are linguistic hubs, with approximately 30 or more languages spoken in each. These results suggest that our embedding can serve as an indicator of a language’s international reach and linguistic diversity, which is not solely reflected by its degree.
@@
Figure 5: Language or country diversity for the top 15 highest-degree nodes. Panel (a) shows violin plots of angular distances between each language and its neighboring countries, with colors indicating script type. Panel (b) presents analogous plots for countries, with colors representing geographic region. In both panels, nodes are ordered in descending order of degree, with each node’s degree (in brackets) indicated next to its label. Only the central 95% of the data is plotted—that is, data between the 2.5th and 97.5th quantiles are shown. A black line highlights the median value in each plot.
@@
In addition, we investigated the human metabolic network, defined as metabolites connected to the reactions they participate in King et al. (2016) and the network of food ingredients based on the flavor compounds they share Ahn et al. (2011) (see Section IV.3). In both cases, B-Mercator is able to reproduce topological properties of this biparite network such as the degree distributions and clustering spectra. See Supplementary Section 6 for more details.
@@
Finally, we applied the bipartite greedy routing protocol to the embeddings derived from these real-world networks. For the Unicodelang network, the highest success rate—based on the $A-A$ BGR variant—is observed at an embedding dimension of $D=4$. However, variations in the success probability ($p_{s}$) across different embedding dimensions are minimal, likely due to the relatively low value of $\beta_{b}$. In contrast, the highest $p_{s}$ for both the Metabolic and Flavor networks is achieved at $D=1$. A summary of these findings is provided in Supplementary Table 2.
II.5 Case study on the graph machine learning tasks
@@
Graph Machine Learning (Graph ML) focuses on extracting patterns, making predictions, and uncovering insights from graph-structured data Hu et al. (2020); Xia et al. (2021). This data is typically defined as a set of entities (nodes) with complex relationships (links), defining a unipartite graph $\mathcal{G}_{n}$. Nodes are enriched with a set of features, defining a bipartite network of nodes and features $\mathcal{G}_{n,f}$. Within the Graph ML community, two key tasks are commonly used to evaluate and rank network embedding models: node classification (NC) and link prediction (LP). In turn, network embeddings can be broadly categorized into supervised and unsupervised approaches. Supervised embeddings, such as those learned by graph neural networks (GNNs), use node labels in the training set to inform the learning process for classification tasks. In contrast, unsupervised methods leverage only the network structure and, optionally, node features to generate low-dimensional representations of the data. These maps can then be used for multiple downstream tasks by integrating additional classification models.
Our embedding method B-Mercator belongs to the class of unsupervised graph embeddings that leverage only the nodes’ feature matrices. This is possible thanks to the findings in Aliakbarisani et al. (2023), which show that nodes and their associated features define a bipartite network, $\mathcal{G}_{n,f}$, with strong geometric properties. We used B-Mercator to find an embedding of the nodes-features bipartite network in the common similarity space, which we subsequently used to perform NC and LP tasks in a supervised manner. To highlight the importance of such embeddings, we compared B-Mercator with D-Mercator Jankowski et al. (2023), which produces multidimensional hyperbolic maps of unipartite networks $\mathcal{G}_{n}$ without using information from the nodes’ features. We selected embedding dimensions $D=1$ and $D=2$ to map the node features into the bipartite-$\mathbb{S}^{1}$ and bipartite-$\mathbb{S}^{2}$ models using B-Mercator, and to map the unipartite network into the $\mathbb{S}^{1}$ and $\mathbb{S}^{2}$ models using D-Mercator.
We also compared our model-driven approach with existing state-of-the-art graph embedding methods in both node classification and link prediction tasks. For further details on the methods used, see Supplementary Section 7. We selected seven graph datasets commonly used in machine learning research, each with varying levels of correlations between the graph, nodes’ features, and nodes’ labels, (see Supplementary Section 8 for details). It has been shown recently that the performance metrics of Graph ML tasks can vary significantly depending on these correlations Jankowski et al. (2024). For instance, in the node classification tasks, adding features can be detrimental when the correlation between node features and network structure is very low. In contrast, adding features significantly enhances the results when the correlation is high. Among the analyzed networks, Cora and Citeseer exhibit a strong correlation between network structure and node features. In contrast, IMDB, Wisconsin, Texas, and Cornell show relatively low correlation, while Film demonstrates almost no correlation (see Supplementary Table S3 for more details).
@@
Figure 6: Case study on machine learning tasks. (a) Accuracy of the node classification task and (b) balanced precision of the distance-based link prediction task for the Wisconsin dataset. The train/test split is 20/80 for nodes in NC and 90/10 for links in LP, where the test set is balanced by randomly adding an equal number of negative links selected from non-existing links. The results are averaged over 10 different splits. Our methods are highlighted with diagonal hatches. The abbreviations of the algorithms are as follows: DW – DeepWalk, R2V – Role2Vec, LE – Laplacian Eigenmaps, FN – FeatherNode. The numeric value in brackets indicates the embedding dimension. All other parameters are set to their default values. The methods are sorted by the median accuracy in NC and balanced precision in LP tasks. In addition, the methods are grouped and colored by input data type, i.e., Network-based methods use only network topology, Feature-based methods use only nodes’ feature matrix, and Feature$+$Network-based methods merge two things to construct the network embedding. (c) Visualization of the bipartite embedding with B-Mercator in $D=1$ for the Wisconsin dataset. We plot the positions of the nodes and color them based on the metadata. (d) Average rank in link prediction and node classification tasks across seven datasets. We plot the medium rank value with the error bars depicting the interquartile range (IQR). The size of the markers is inversely proportional to the embedding dimension and their shapes correspond to the input data type. Our methods are outlined with a thicker marker border.
@@
In Fig. 6b, we show the precision of the different methods. In this case, methods using only the network topology, or a combination of network structure and node features, deliver the highest precisions. For this task, our embeddings with D-Mercator ($D=1$ and $D=2$), using only the network topology, are competitive and achieve a performance similar to MUSAE in dimension $D=8$, demonstrating the versatility of low-dimensional embeddings derived from our approach. Moreover, among feature-based methods, B-Mercator in $D=1$ and $D=2$ outperforms all other approaches. This highlights the adaptability of our methods to different tasks depending on the underlying data representation and task requirements. Detailed results for additional datasets can be found in Supplementary Figures S23–S46.
Figure 6d provides a summary of the results across multiple datasets. For each dataset, the rank of each method is calculated for both NC and LP tasks, and the median rank is plotted along with the interquartile range (IQR). B-Mercator in low dimensions ($D=1$ and $D=2$) consistently ranks among the top approaches for the NC task and outperforms all feature-based methods for the LP task, for which D-Mercator is the best method in dimension $D=1$. These results demonstrate the reliability and adaptability of our methods across datasets. The marker size in the plot, which inversely represents the embedding dimension, further underscores the efficiency of our low-dimensional embeddings compared to high-dimensional alternatives. In Supplementary Figures S30 and S38, we provide a more detailed view of the rankings for the NC and LP tasks.
III Discussion
@@
Several model-based embedding methods have already proved effective for unipartite networks, yet comparable solutions for bipartite networks have remained underexplored. In this work, we addressed this gap by introducing B-Mercator, a novel geometric model-based embedding algorithm specifically designed for bipartite networks. By mapping bipartite structures into hyperbolic space, B-Mercator offers a powerful and interpretable way to capture community structure, hierarchical organization, and topological relationships.
To demonstrate the versatility of B-Mercator, we applied it to embed real-world bipartite systems. The analysis of the language network (countries–spoken languages), the metabolic network (metabolites–reactions), and the Flavor network (ingredients–chemical compounds) show that the embeddings not only correlate well with metadata but also retain the essential characteristics of each dataset. Furthermore, we evaluated B-Mercator on several node classification and link prediction tasks. It consistently outperformed all unsupervised methods for node classification and emerged as the best performer among feature-based embeddings for link prediction.
We stress an important advantage of our model-based embeddings. The hyperbolic maps can be used to generate synthetic networks with node features that closely resemble their real-world counterparts (see Supplementary Figures 46-48). By doing so, we safeguard any sensitive information derived from real complex networks, such as personal connections, transactional data, or proprietary interactions. Thus, we enable the secure sharing of structural data without compromising the integrity of the original network or revealing sensitive information.
@@
IV Methods
IV.1 Bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model
In the bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model – an extension of the bipartite-$\mathbb{S}^{1}/\mathbb{H}^{2}$ Aliakbarisani et al. (2023) – we assign to each node (of type A or B) a hidden degree ($\kappa_{A}$ or $\kappa_{B}$) and the position in the $D$-dimensional similarity space chosen uniformly at random, and represented as a point on a $D$-dimensional sphere. Each node of type A and B is assigned a vector $\mathbf{x}_{i}\in\mathbb{R}^{D+1}$ with $||\mathbf{x}_{i}||=R$. For instance, when $D=1$ the similarity space is represented as a circle, whereas for $D=2$ it is a sphere (see Figure 1).
The connection probability between node $u$ of type A and node $v$ of type B takes the form of gravity law:
@@ IV.1 Bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model
The number of nodes of type A (type B) is $N_{A}$($N_{B}$), for convenience and without loss of generality, we set the density of nodes of type A in the $D$-sphere to one so that
@@
The separation $\Delta\theta_{uv}=\arccos\left(\frac{\mathbf{x}_{u}.\mathbf{x}_{v}}{R^{2}}\right)$ represents the angular distance between nodes $u$ and $v$ in the $D$-dimensional similarity space. The parameter $\beta_{b}$ (with $\beta_{b}>D$) controls the coupling between the resulting topology and the underlying metric space. Lastly, the parameter $\mu_{b}$ controls the average degree of nodes of type A and is defined as
@@
whereas the average degree of type B nodes is set by $\langle k_{B}\rangle=\frac{N_{A}}{N_{B}}\langle k_{A}\rangle$. By choosing distributions for the hidden degrees $\kappa_{A}$ and $\kappa_{B}$ and inverse temperature $\beta_{b}$, we can generate bipartite networks with any desired degree distributions and varying levels of geometric properties.
Interestingly, the bipartite-$\mathbb{S}^{D}$ model can be represented in purely geometric terms as the bipartite-$\mathbb{H}^{D+1}$ model. This is achieved by mapping the hidden degrees of each type A and B nodes to radial coordinates while preserving their positions on the sphere $\mathbb{S}^{D}$. Specifically, the transformation for type A nodes has the form (similarly for type B nodes)
@@
where ${\kappa}_{u,0}$(${\kappa}_{v,0}$) is the smallest hidden degree for type A (type B) nodes. We can rewrite Eq. (1) as
@@
With this transformation, the space represented by the radial position of each node, along with its angular position on the sphere, becomes the hyperboloid model of the hyperbolic space of dimension $D+1$. Consequently, the connection probability between nodes $u$ and $v$ becomes a function of $x_{uv}$, which is a good approximation of the hyperbolic distance between them.
IV.2 B-Mercator in details
We adopt the code of $D$-Mercator Jankowski et al. (2023) for the embedding bipartite networks. Here, we provide the overview of the differences between the embeddings for the unipartite and bipartite networks.

### v2

@@ I.4 Embedding of real bipartite networks
Figure 4: Visualization of the bipartite-$\mathbb{S}^{1}$ embedding of the Unicodelang dataset per country or language. Panels (a, b, c) show countries where a given language is spoken, i.e., the neighbors of the language node. The size of the nodes is proportional to the number of language speakers in that country. The color corresponds to the geographical region in which the country is located. A star marker indicates the position of a given language. In panels (d, e, f), we depict all languages spoken in a given country, i.e., the neighbors of the country node. The size of the nodes is proportional to the fraction of speakers of a given language. The color represents that language’s script. A cross marker indicates the position of a given country.
Using B-Mercator, we embedded the Unicodelang dataset in various dimensions. The inferred embeddings are able to reproduce the topological properties of the network (see Supplementary Figure 17). In Figure 4, we show a dual embedding representation of this dataset in dimension $D=1$. First, we focus on the countries in which a given language is used. For instance, in Figure 4a, we plot all countries in which English is spoken, i.e., the neighbors of the English language in the bipartite network. One can notice that English is located closer to India or the Philippines than to the United States of America (USA). This can be explained by the fact that many different languages are spoken in the USA, which influence its position in the bipartite map. Indeed, in Figure 4b, where we plot all neighbors of the Spanish language, the USA is located close to the Spanish language. Lastly, in Figure 4c one can observe that countries from the French colonialism are concentrated in the similarity space and lie close to the French language, in contrast to France and other European countries. We can also shift the perspective: instead of examining the neighbors of each language, we can analyze the neighbors of a given country. In the bottom row of Figure 4, we plot the language neighbors of India, China, and Brazil. Countries are often located in the embedding space close to the most widely spoken language. Additional examples are provided in Supplementary Figure 15, where we explore the neighbors of languages such as Hindi, Swahili, Catalan, Persian, Korean, Dutch, Russian, and Arabic. Similarly, in Supplementary Figure 16, we depict the neighbors of countries including Canada, Turkey, Indonesia, Cameroon, Tanzania, Greece, the Philippines, and Bolivia.
@@
Figure 5: Language or country diversity for the top 15 highest-degree nodes. Panel (a) shows violin plots of angular distances between each language and its neighboring countries, with colors indicating script type. Panel (b) presents analogous plots for countries, with colors representing geographic region. In both panels, nodes are ordered in descending order of degree, with each node’s degree (in brackets) indicated next to its label. Only the central 95% of the data is plotted–that is, data between the 2.5th and 97.5th quantiles are shown. A black line highlights the median value in each plot.
The hyperbolic bipartite embedding enables us to examine the concentration of countries sharing the same language. We select the top 15 languages with the highest degree and compute the angular distance from each language to its neighboring countries. A small average angular distance may indicate that these countries are highly similar, whereas a broad distribution suggests a more international language. Figure 5a displays the diversity for these high-degree languages. As expected, the angular distance distribution for English is broad, indicating connections with countries distributed throughout the similarity space. Instead, the angular distances for Fula are relatively small. Fula, a Senegambian language spoken primarily in West and Central Africa, is concentrated in a specific region of our embedding space. Similarly, we can examine the linguistic diversity within each country. Following the previous approach, we selected the top 15 countries with the highest degree and computed the angular distances to each neighboring language. In a given country, if the languages are more similar, their angular positions in the similarity space should be more concentrated. Figure 5b shows that India and Canada exhibit a broader distribution of angular distances, reflecting the presence of a diverse array of languages. In contrast, Russia and Cameroon display a narrower distribution, indicating a more homogeneous set of languages. Notably, all these countries are linguistic hubs, with approximately 30 or more languages spoken in each. These results suggest that our embedding can serve as an indicator of a language’s international reach and linguistic diversity, which is not solely reflected by its degree.
@@
In addition, we investigated the human metabolic network, defined as metabolites connected to the reactions they participate in King et al. (2016) and the network of food ingredients based on the flavor compounds they share Ahn et al. (2011) (see Section III.3). In both cases, B-Mercator is able to reproduce topological properties of this bipartite network such as the degree distributions and clustering spectra. See Supplementary Note 6 for more details.
@@
Finally, we applied the bipartite greedy routing protocol to the embeddings derived from these real-world networks. For the Unicodelang network, the highest success rate—based on the AAA-A BGR variant—is observed at an embedding dimension of $D=4$. However, variations in the success probability ($p_{s}$) across different embedding dimensions are minimal, likely due to the relatively low value of $\beta_{b}$. In contrast, the highest $p_{s}$ for both the Metabolic and Flavor networks is achieved at $D=1$. A summary of these findings is provided in Supplementary Tables S2–S5.
I.5 Case study on the graph machine learning tasks
@@
Graph Machine Learning (Graph ML) focuses on extracting patterns, making predictions, and uncovering insights from graph-structured data Hu et al. (2020); Xia et al. (2021). This data is typically defined as a set of entities (nodes) with complex relationships (links), defining a unipartite graph $\mathcal{G}_{n}$. Nodes are classified in different categories (labels) and are enriched with a set of features, defining a bipartite network of nodes and features $\mathcal{G}_{n,f}$. Within the Graph ML community, two key tasks are commonly used to evaluate and rank network embedding models: node classification (NC) and link prediction (LP). In turn, network embeddings can be broadly categorized into supervised and unsupervised approaches. Supervised embeddings, such as those learned by graph neural networks (GNNs), use node labels in the training set to inform the learning process for classification tasks. In contrast, unsupervised methods leverage only the network structure and, optionally, node features to generate low-dimensional representations of the data. These maps can then be used for multiple downstream tasks by integrating additional classification models.
@@
Figure 6: Case study on machine learning tasks. (a) Accuracy of the node classification (NC) task and (b) balanced precision of the distance-based link prediction (LP) task for the Wisconsin dataset. The train/test split is 20/80 for nodes in NC and 90/10 for links in LP, where the test set is balanced by randomly adding an equal number of negative links selected from non-existing links. The results are averaged over 10 different splits. Our methods are highlighted with diagonal hatches. The abbreviations of the algorithms are as follows: DW – DeepWalk, R2V – Role2Vec, LE – Laplacian Eigenmaps, FN – FeatherNode. With a star, we highlighted methods that use both network topology and node features for embedding. The numeric value in brackets indicates the embedding dimension. All other parameters are set to their default values. The methods are sorted by the median accuracy in NC and balanced precision in LP tasks. In addition, the methods are grouped and colored by input data type, i.e., Network-based methods use only network topology, Feature-based methods use only nodes’ feature matrix, and Feature++Network-based methods merge two things to construct the network embedding. (c) Visualization of the bipartite embedding with B-Mercator in $D=1$ for the Wisconsin dataset. We plot the positions of the nodes and color them based on the metadata. (d) Average rank in link prediction and node classification tasks across seven datasets. We plot the median rank value with the error bars depicting the interquartile range (IQR), i.e., the interval between the first and third quartiles that captures the middle 50% of the rank distribution. The size of the markers is inversely proportional to the embedding dimension and their shapes correspond to the input data type. Our methods are outlined with a thicker marker border.
Our embedding method B-Mercator belongs to the class of unsupervised graph embeddings that leverage only the nodes’ feature matrices. This is possible thanks to the findings in Aliakbarisani et al. (2025), which show that nodes and their associated features define a bipartite network, $\mathcal{G}_{n,f}$, with strong geometric properties. We used B-Mercator to find an embedding of the nodes-features bipartite network in the common similarity space, which we subsequently used to perform NC and LP tasks in a supervised manner. To highlight the importance of such embeddings, we compared B-Mercator with D-Mercator Jankowski et al. (2023), which produces multidimensional hyperbolic maps of unipartite networks $\mathcal{G}_{n}$ without using information from the nodes’ features. It has been shown that real complex networks can be accurately embedded in low dimensional hyperbolic spaces Jankowski et al. (2024). Thus, in this work we selected embedding dimensions $D=1$ and $D=2$ to map the node features into the bipartite-$\mathbb{S}^{1}$ and bipartite-$\mathbb{S}^{2}$ models using B-Mercator, and to map the unipartite network into the $\mathbb{S}^{1}$ and $\mathbb{S}^{2}$ models using D-Mercator.
We also compared our model-driven approach with existing state-of-the-art graph embedding methods in both node classification and link prediction tasks. Prior work shows that efficient encoding can often be achievable in low-dimensional spaces Gu et al. (2021). Therefore, we consider embedding dimensions between 2 and 16. For further details on the methods used, see Supplementary Note 7. We selected seven graph datasets commonly used in machine learning research, each with varying levels of correlations between the graph, nodes’ features, and nodes’ labels (see Supplementary Note 8 for details). It has been shown recently that the performance metrics of Graph ML tasks can vary significantly depending on these correlations Jankowski et al. (2024). For instance, in the node classification tasks, adding features can be detrimental when the correlation between node features and network structure is very low. In contrast, adding features significantly enhances the results when the correlation is high. Among the analyzed networks, Cora and Citeseer exhibit a strong correlation between network structure and node features. In contrast, IMDB, Wisconsin, Texas, and Cornell show relatively low correlation, while Film demonstrates almost no correlation (see Supplementary Table S6 for more details).
@@
In Fig. 6b, we show the precision of the different methods. In this case, methods using only the network topology, or a combination of network structure and node features, deliver the highest precisions. For this task, our embeddings with D-Mercator ($D=1$ and $D=2$), using only the network topology, are competitive and achieve a performance similar to MUSAE in dimension $D=8$, demonstrating the versatility of low-dimensional embeddings derived from our approach. Moreover, among feature-based methods, B-Mercator in $D=1$ and $D=2$ outperforms all other approaches. This highlights the adaptability of our methods to different tasks depending on the underlying data representation and task requirements. Detailed results for additional datasets can be found in Supplementary Figures S25–S50.
Figure 6d provides a summary of the results across multiple datasets. For each dataset, the rank of each method is calculated for both NC and LP tasks, and the median rank is plotted along with the interquartile range (IQR). B-Mercator in low dimensions ($D=1$ and $D=2$) consistently ranks among the top approaches for the NC task and outperforms all feature-based methods for the LP task, for which D-Mercator is the best method in dimension $D=1$. These results demonstrate the reliability and adaptability of our methods across datasets. The marker size in the plot, which inversely represents the embedding dimension, further underscores the efficiency of our low-dimensional embeddings compared to high-dimensional alternatives. In Supplementary Figures S32 and S42, we provide a more detailed view of the rankings for the NC and LP tasks.
II Discussion
@@
Several model-based embedding methods have already proved effective for unipartite networks, yet comparable solutions for bipartite networks have remained underexplored. In this work, we addressed this gap by introducing B-Mercator, a geometric model-based embedding algorithm specifically designed for bipartite networks. By mapping bipartite structures into hyperbolic space, B-Mercator offers a powerful and interpretable way to capture community structure, hierarchical organization, and topological relationships.
To demonstrate the versatility of B-Mercator, we applied it to embed real-world bipartite systems. The analysis of the language network (countries–spoken languages), the metabolic network (metabolites–reactions), and the Flavor network (ingredients–chemical compounds) show that the embeddings not only correlate well with metadata but also retain the essential characteristics of each dataset. Furthermore, we evaluated B-Mercator on several datasets on the node classification and link prediction tasks. It consistently outperformed all unsupervised methods for node classification and emerged as the best performer among feature-based embeddings for link prediction.
We stress an important advantage of our model-based embeddings. The hyperbolic maps can be used to generate synthetic networks with node features that closely resemble their real-world counterparts (see Supplementary Figures 51-53). By doing so, we safeguard any sensitive information derived from real complex networks, such as personal connections, transactional data, or proprietary interactions. Thus, we enable the secure sharing of structural data without compromising the integrity of the original network or revealing sensitive information.
@@
III Methods
III.1 Bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model
In the bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model – an extension of the bipartite-$\mathbb{S}^{1}/\mathbb{H}^{2}$ Aliakbarisani et al. (2025) – we assign to each node (of type-A or type-B) a hidden degree ($\kappa_{A}$ or $\kappa_{B}$) and the position in the $D$-dimensional similarity space chosen uniformly at random, and represented as a point on a $D$-dimensional sphere. Each node of type-A and -B is assigned a vector $\mathbf{x}_{i}\in\mathbb{R}^{D+1}$ with $\mathbf{x}_{i}||=R$. For instance, when $D=1$ the similarity space is represented as a circle, whereas for $D=2$ it is a sphere (see Figure 1).
The connection probability between type-A node $u$ and type-B node $v$ takes the form of gravity law:
@@ III.1 Bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ model
The number of type-A (type-B) nodes is $N_{A}$($N_{B}$), for convenience and without loss of generality, we set the density of type-A nodes in the $D$-sphere to one so that
@@
The separation $\Delta\theta_{uv}=\arccos\left(\frac{\mathbf{x}_{u}.\mathbf{x}_{v}}{R^{2}}\right)$ represents the angular distance between nodes $u$ and $v$ in the $D$-dimensional similarity space. The parameter $\beta_{b}$ (with $\beta_{b}>D$) controls the coupling between the resulting topology and the underlying metric space. Lastly, the parameter $\mu_{b}$ controls the average degree of type-A nodes and is defined as
@@
whereas the average degree of type-B nodes is set by $\langle k_{B}\rangle=\frac{N_{A}}{N_{B}}\langle k_{A}\rangle$. By choosing distributions for the hidden degrees $\kappa_{A}$ and $\kappa_{B}$ and inverse temperature $\beta_{b}$, we can generate bipartite networks with any desired degree distributions and varying levels of geometric properties.
The bipartite-$\mathbb{S}^{D}$ model can be represented in purely geometric terms as the bipartite-$\mathbb{H}^{D+1}$ model. This is achieved by mapping the hidden degrees of each type-A and -B nodes to radial coordinates while preserving their positions on the $D$-sphere. Specifically, the transformation for type-A nodes has the form (similarly for type-B nodes)
@@
where ${\kappa}_{u,0}$(${\kappa}_{v,0}$) is the smallest hidden degree for type-A (type-B) nodes. We can rewrite Eq. (1) as
@@
With this transformation, the space represented by the radial position of each node, along with its angular position on the sphere, becomes the native representation of the hyperbolic space of dimension $D+1$. Consequently, the connection probability between nodes $u$ and $v$ becomes a function of $x_{uv}$, which is a good approximation of the hyperbolic distance between them Serrano and Boguñá (2022).
III.2 B-Mercator in detail
We adopt the code of $D$-Mercator Jankowski et al. (2023) for embedding bipartite networks. Here, we provide an overview of the differences between the embeddings for the unipartite and bipartite networks. See also Supplementary Note 12 for more technical details.